# Spin Relaxation Mechanisms and Nuclear Spin Entanglement of the V Center in hBN

### v1

@@ 2.2 Effect of Nuclear Spin Entanglement on
In this section, Model 4 is used to simulate the spin‐relaxation dynamics at G, which is the same strength performed in the experiment[22]. Figs. 4(a) - (c) depict the variations in a population of polarization among a single electron spin, one of the three first-nearest nuclear spins, and the sum of two distant bath spins, respectively. These results indicate that the electron spin gradually transfers its polarization to all surrounding nuclei, including both nearest neighbors and the remote bath spins, as evidenced by the sharp increase in nuclear‐spin polarization in Figs. 4(b) and (c). As time progresses, the electron‐spin polarization continues to decay until full depolarization is achieved; simultaneously, the nearest‐neighbor nuclei also relax back toward depolarization. In contrast, once the two distant bath nuclei acquire polarization from the electron spin, they subsequently begin to transfer polarization back to the first‐nearest nitrogen nuclei. Finally, all spin populations will converge towards a value of 0.5, indicating that the spins are depolarizing and approaching thermal equilibrium. This coherent oscillation of the first nearest nitrogen nuclei, as exhibited in Fig. 4(d), can be explained by the oscillation between and states. Such the polarization population changes are following the spin flip-flop processes, as explained earlier.
Now, turning to quantify the time, Fig. 4(a) showcases that the simulated polarization did not reach the thermal equilibrium point as the simulation has reached the allowed runtime in our supercomputer, which is 20 days. We anticipated that the simulation could take up to 80 days to reach the equilibrium point. To resolve this limitation, we simulated the polarization dynamics until it reached a maximum of 20 running days. Then we fitted the curve and extrapolated the rest of the behavior. This is still valid given that the polarization will, in principle, reach the equilibrium point at the polarization equal to 0.5. Furthermore, we have cross-checked with the polarization behavior at = 800 G as shown in Fig. 3(d) and found the equilibrium pattern. Note that the polarization can be simulated faster in the field close to GSLAC due to the narrower gap between and , as shown in Fig. S1 in Supplementary S1. As a result, the fitted value of at = 90 G yields 694.77 s, which is quite consistent with the experimental value of 1 ms reported at 10 K in h10B15N[22]. For the small discrepancy, we hypothesize that this might arise from the absence of some correlations between different clusters, which are not captured even by our six-spin cluster approach.
@@ 2.3 Dependence on Magnetic Field
In low fields, decreases as approaches GSLAC near 1250 G. This can be attributed to the narrowing of the Zeeman splitting between and levels, as displayed in Fig. S1 in Supplementary S1. The reduced energy gap enhances electron–nuclear flip–flop transitions mediated by the non‐secular hyperfine term , while the dipolar interaction term promotes faster nuclear–nuclear spin flip–flops by increasing bath fluctuations, as evidenced by the larger population changes of bath spins in the higher field shown in Supplementary S2. As a result, both mechanisms accelerate decoherence processes, leading to a shorter time with higher B strength. It is important to note that at lower fields ( G), the values presented in Fig. 5(a) are likely overestimated. This is because the current model reduces the Hilbert space by applying the reduced density matrix formalism to consider only the and sublevels. In practice, spin relaxation pathways involving flip-flop processes between and are also possible and should be taken into account. Accordingly, the total should be given by the relation
@@
where and denote the relaxation times associated with spin flip-flop transitions between and , and between and , respectively. To address this limitation based on the same reduce density matrix, we computed the time under a field applied in the opposite direction. As shown in Fig. 5, the total is slightly lower than due to the presence of additional spin relaxation pathways. However, beyond a strength of 400 G, the total converges to values comparable to . This behavior arises because, at higher fields, the energy gap between and narrows while the gap between and widens, as illustrated in Supplementary Fig. S1. Consequently, spin relaxation via the channel dominates, and the overall relaxation rate becomes effectively determined by this single pathway.
@@
In the GSLAC region, the system exhibits different relaxation dynamics. Here, the and eigenstates undergo strong hybridization with nearby nuclear spin levels, resulting in oscillations in the population dynamics rather than the monotonic exponential decay, as shown in Supplementary S2. Under these conditions, a time cannot be extracted in the conventional sense based on an exponential fitting. This leads to an abrupt discontinuity in the 1/ behavior. This behavior has also been observed in 13C in NV center in diamond[42], as well as 13C and 29Si in SiC[38].
@@
In high fields, although nuclear–nuclear dipolar interactions persist and continue to contribute to bath-induced noise, their influence diminishes with increasing , as indicated by the reduced population fluctuations of bath spins in Supplementary S2. Furthermore, the increased energy separation between the and states suppresses hyperfine-mediated flip–flop transitions. Consequently, the overall relaxation rate decreases, leading to an enhancement of the time in this regime. It is important to note, however, that the computed values in this regime may still be slightly overestimated. This overestimation arises from the fact that, similar to the situation in low fields, the current model considers relaxation via the channel while neglecting potential relaxation through the pathway. As shown in Fig. S1 in Supplementary S1, the Zeeman splitting between the and states increases substantially with . At sufficiently high fields, the energy gaps between and both states become approximately equal, making relaxation through both pathways equally probable. In this limit, the correct should be halved relative to the value obtained from a single-channel calculation; that is, .
@@
For the field beyond 2000 G, this aspect remains a computational limitation in the present work, as indicated in Supplementary S2. This is due to the very large Zeeman splitting of and at high fields; thus, capturing the oscillation over such a wide energy range requires significantly increased computational resources. Addressing this in future studies would further refine the accuracy of predictions in the higher field regime.
@@ 3 Discussion
Overall, this work has established an effective and reliable model for describing relaxation dynamics in V centers in hBN. The insights gained into the underlying decoherence mechanisms provide a foundation for further exploration of spin dynamics in solid-state quantum systems. For instance, future studies could extend this framework to include additional relaxation mechanisms, such as interactions with nearby paramagnetic electron spins or boron nuclei, and investigate their contributions to the overall spin relaxation behavior. Moreover, this understanding could inform strategies to engineer longer times, either through defect environment control or external manipulation. Such advancements would be valuable for optimizing V centers for applications in quantum sensing and quantum memory technologies.
@@ 4 Methods
To solve the spin dynamics governed by the Lindblad master equation [Eq. (6)], we use our in-house MPI-parallelized C code implementing a fourth-order Runge-Kutta integration scheme. A timestep of 2.5 s is employed, as validated in Supplementary S5. To simulate the dynamics of V centers, we incorperate system-specific parameters, including the ZFS and HF interaction tensors. The ZFS is obtained from real-space calculations, which include all nuclei within a 30-Å radius from the V defect in bulk hBN to resolve the finite size effect[44]. This corresponds to a total of 12,702 atoms in the supercell. The HF interaction is set to 3.479 GHz, as calculated using DFT together with the spin decotamination approach described in Ref.[24]. The spin relaxation is then evaluated by monitoring the population of the V electron spin as it evolves toward thermal equilibrium. Here, the initial density matrix is prepared such that the V electron spin is polarized with 70% population in the state and 30% in the state, while all nuclear spins are initialized in the thermal equilibrium depolarized states.
@@ Acknowledgments
CC initiated the study and carried out the simulations. The manuscript was written by CC with inputs from TV and VI. The work was supervised with TV and VI.

### v2

@@ 2.6 Effect of Nuclear Spin Entanglement on
In this section, Model 4 is used to simulate the spin‐relaxation dynamics at G, which is the same strength used in the experiment[10.1038/s41467-023-44494-3]. Figs. 5(a) - (c) depict the variations in a population of polarization among a single electron spin, one of the three first-nearest nuclear spins, and the sum of two distant bath spins, respectively. These results indicate that the electron spin gradually transfers its polarization to all surrounding nuclei, including both nearest neighbors and the remote bath spins, as evidenced by the sharp increase in nuclear‐spin polarization in Figs. 5(b) and (c). As time progresses, the electron‐spin polarization continues to decay until full depolarization is achieved. Both the nearest-neighbor and the more distant nuclear spins exhibit an initial transient period (30 s) during which the nuclear spin polarization rises to positive values before decaying. After this transient regime, the entire bath enters a steady-state period, where similar dynamics are preserved over longer timescales, leading to a gradual buildup of polarization in the bath. At a later time, not achieved in our simulations, all spin populations will converge towards a value of 0.5, indicating that the spins are depolarizing and approaching thermal equilibrium. The fast coherent oscillation of the first nearest nitrogen nuclei, as exhibited in Fig. 5(d), can be explained by the oscillation between the and the states induced by the hyperfine interaction.
Now, turning to quantify the time, Fig. 5(a) showcases that the simulated polarization did not reach the thermal equilibrium point as the simulation has reached the allowed runtime in our supercomputer, which is 20 days. We anticipated that the simulation could take up to 80 days to reach the equilibrium point. To resolve this limitation, we simulated the polarization dynamics until it reached a maximum of 20 running days. Then we fitted the curve and extrapolated the rest of the behavior. This is still valid given that the polarization will, in principle, reach the equilibrium point at the polarization equal to 0.5. Furthermore, we have cross-checked with the polarization behavior at = 800 G as shown in Fig. 4(d) and found the equilibrium pattern. Note that the polarization can be simulated faster in the field close to GSLAC due to the narrower gap between and , as shown in Fig. 1. As a result, the fitted value of at = 90 G yields 694.77 s, which is quite consistent with the experimental value of 1 ms reported at 10 K in h10B15N[10.1038/s41467-023-44494-3]. For the small discrepancy, we hypothesize that this might arise from the absence of some correlations between different clusters, which are not captured even by our six-spin cluster approach. Moreover, while this work employed a ZFS value of 3.479 GHz, the ZFS increases to approximately 3.63 GHz at low temperatures [10.1038/s41467-021-24725-1]. The increased ZFS enhances the energy separation between the three states in the interval. Consequently, the time is expected to be slightly larger at low magnetic fields, bringing our results into closer agreement with the experiment.
@@ 2.7 Dependence on Magnetic Field
In low fields, decreases as approaches GSLAC near 1250 G. This can be attributed to the narrowing of the Zeeman splitting between and levels, as displayed in Fig. 1. The reduced energy gap enhances electron–nuclear flip–flop transitions mediated by the non‐secular hyperfine term , while the dipolar interaction term promotes faster nuclear–nuclear spin flip–flops by increasing bath fluctuations, as evidenced by the larger population changes of bath spins in the higher field shown in Supplementary S1. As a result, both mechanisms accelerate decoherence processes, leading to a shorter time with higher B strength. It is important to note that at lower fields ( G), the values presented in Fig. 6(a) are likely overestimated. This is because the current model reduces the Hilbert space by applying the reduced density matrix formalism to consider only the and sublevels. In practice, spin relaxation pathways involving flip-flop processes between and are also possible and should be taken into account. Accordingly, the total should be given by the relation
@@
where and denote the relaxation times associated with spin flip-flop transitions between and , and between and , respectively. To address this limitation based on the same reduce density matrix, we computed the time under a field applied in the opposite direction. As shown in Fig. 6, the total is slightly lower than due to the presence of additional spin relaxation pathways. However, beyond a strength of 400 G, the total converges to values comparable to . This behavior arises because, at higher fields, the energy gap between and narrows while the gap between and widens, as illustrated in Fig. 1. Consequently, spin relaxation via the channel dominates, and the overall relaxation rate becomes effectively determined by this single pathway.
@@
In the GSLAC region, the system exhibits different relaxation dynamics. Here, the and eigenstates undergo strong hybridization with nearby nuclear spin levels, resulting in oscillations in the population dynamics rather than the monotonic exponential decay, as shown in Supplementary S1. Under these conditions, a time cannot be extracted in the conventional sense based on an exponential fitting. This leads to an abrupt discontinuity in the 1/ behavior. This behavior has also been observed in 13C in NV center in diamond[10.1103/PhysRevB.101.155203], as well as 13C and 29Si in SiC[10.1038/s41524-021-00673-8].
@@
In high fields, although nuclear–nuclear dipolar interactions persist and continue to contribute to bath-induced noise, their influence diminishes with increasing , as indicated by the reduced population fluctuations of bath spins in Supplementary S1. Furthermore, the increased energy separation between the and states suppresses hyperfine-mediated flip–flop transitions. Consequently, the overall relaxation rate decreases, leading to an enhancement of the time in this regime. It is important to note, however, that the computed values in this regime may still be slightly overestimated. This overestimation arises from the fact that, similar to the situation in low fields, the current model considers relaxation via the channel while neglecting potential relaxation through the pathway. As shown in Fig. 1, the Zeeman splitting between the and states increases substantially with . At sufficiently high fields, the energy gaps between and both states become approximately equal, making relaxation through both pathways equally probable. In this limit, the correct should be halved relative to the value obtained from a single-channel calculation; that is, .
@@
For the field beyond 2000 G, this aspect remains a computational limitation in the present work, as indicated in Supplementary S1. This is due to the very large Zeeman splitting of and at high fields; thus, capturing the oscillation over such a wide energy range requires significantly increased computational resources. Addressing this in future studies would further refine the accuracy of predictions in the higher field regime.
@@ 3 Discussion
While this work investigates the time in bulk hBN, dimensionality of the sample can also influence the results. In Model 4 of bulk hBN, the two distant nitrogen nuclei can sit either in the defected layer or in the neighboring layers. In contrast, in a monolayer hBN geometry, the two distant nitrogen nuclei would necessarily lie within the same plane and may therefore be located farther from the vacancy than in the bulk case. This would reduce their hyperfine interaction strength and, consequently, their contribution to decoherence. As a result, the nuclear spin bath-originated time in monolayer hBN is expected to be longer than in bulk or multilayer systems.
Overall, this work has established an effective and reliable model for describing relaxation dynamics in V centers in hBN. The insights gained into the underlying decoherence mechanisms provide a foundation for further exploration of spin dynamics in solid-state quantum systems. For instance, future studies could extend this framework to include additional relaxation mechanisms, such as interactions with nearby paramagnetic electron spins or boron nuclei, and investigate their contributions to the overall spin relaxation behavior. However, simply replacing the two distant 15N nuclei with 11B nuclei would increase the Hilbert space from 64 to 256, making such simulations computationally demanding with the current code. Improving computational efficiency will therefore be an essential step toward including boron nuclei in a systematic way. Additionally, the work may be continued by explicitly quantifying entanglement in these strongly coupled systems. A deep understanding of spin relaxation may lead to new strategies to engineer longer times, either through defect environment control or external manipulation. Such advancements would be valuable for optimizing V centers for applications in quantum sensing and quantum memory technologies.
@@ 4 Methods
4.1 Computational details
To solve the spin dynamics governed by the Lindblad master equation [Eq. (6)], we use our in-house MPI-parallelized C code implementing a fourth-order Runge-Kutta integration scheme. A timestep of 2.5 s is employed, as validated in Supplementary S4. To simulate the dynamics of V centers, we incorperate system-specific parameters, including the ZFS and HF interaction tensors. The ZFS is set to 3.479 GHz, as calculated using DFT together with the spin decotamination approach described in Ref.[10.1038/s41524-020-0305-x]. The HF interaction is obtained from real-space calculations, which include all nuclei within a 30-Å radius from the V defect in bulk hBN to resolve the finite size effect[10.1038/s42005-024-01668-9]. This corresponds to a total of 12,702 atoms in the supercell. The spin relaxation is then evaluated by monitoring the population of the V electron spin as it evolves toward thermal equilibrium. Here, the initial density matrix is prepared such that the V electron spin is polarized with 70% population in the state and 30% in the state, while all nuclear spins are initialized in the thermal equilibrium depolarized states.
@@ Acknowledgments
C.C. initiated the study and carried out the simulations. The manuscript was written by C.C. with inputs from T.V. and V.I. The work was supervised with T.V. and V.I.